# False Alarms, Real Damage: Adversarial Attacks Using LLM-based Models on Text-based Cyber Threat Intelligence Systems

### v1

@@ 5.2 Target models
Two models are targeted in the experiments. The first is a state-of-the-art binary classifier developed by Dionisio et al. [14] based on Convolutional Neural Networks (CNNs), employed within CTI extraction pipelines to recognize whether a tweet is related to cybersecurity or not. This classifier was selected based on an evaluation [40], which identified it as an effective classifier. Additionally, it was trained on real tweets, which makes it a suitable and realistic target for adversarial evaluation. The second target model is ChatGPT-4o, selected for exceptional performance in classifying security-related and non-security texts [41]. Unlike traditional classifiers, ChatGPT-4o integrates contextual understanding with classification tasks to effectively detect subtle cyber threat patterns. Moreover, the model’s widespread adoption and versatility make it a practical benchmark for testing adversarial robustness in real-world scenarios. By targeting ChatGPT-4o, we aim to evaluate the resilience of cutting-edge language models against evasion attacks. The evasion attack targets both models, while the flooding and poisoning attacks are specific to the specialized classifier [14].
@@ 5.3 Adversarial text generation
ChatGPT-4o was leveraged to generate texts that mimic the structure and terminology of cybersecurity texts while remaining unrelated to real cybersecurity. While LLM-based text generation has limitations, such as potential biases, repetition, and inaccuracies [47], the method proposed in this study exceeded the expectations. To generate effective adversarial texts with ChatGPT-4o, prompts were designed using real tweets. Figure A1 in Appendix A shows the prompt utilized for generating FaN texts, which incorporates key tokens to achieve the desired style while excluding specific security details [48]. An attention-based mechanism is proposed to extract the contextually key tokens from the real cybersecurity tweets. In the following, we first define the problem of generating adversarial text and then explain the generation process and its components as shown in Figure 3.
@@
Figure 3 illustrates the FaN text generation methodology [^2] Code will be available on https://github.com/samanehshf/Fake-CTIs. The attention mechanism identifies the top three tokens, and ChatGPT-4o modifies surrounding words as suggested in the prompt to create adversarial texts that still appear cybersecurity-related. The prompt follows an iterative refinement, where each cycle evaluates the generated FaNs based on their ability to deceive the pre-trained text classifier used in our pipeline. The prompt is considered adequate if the classifier produces a number of FPs (#$\#FP$) exceeding a predefined threshold $\vartheta$. Otherwise, another refinement iteration is executed. The finalized prompt was effective enough to surpass the $\vartheta$ setting of 80% of the input texts. It was then used to generate the adversarial FaN texts, forming the basis for subsequent evaluation.
Human evaluation. The last iteration of prompt optimization involved a human evaluation step, performed to further validate the effectiveness of the generated FaN texts. The data was distributed to three analysts who were informed that the texts were fake and were tasked to identify those that resembled cybersecurity texts. This result was vital to achieving a final prompt surpassing the $\vartheta$ threshold. The importance of the iterative prompt refinement with human validation is demonstrated by the difference between the initial and final FPR presented in Section 6.
@@ 5.4 Evasion attack
The evasion attack was conducted by passing the adversarially generated FaN texts through the target classifiers. A successful attack occurs when a FaN is incorrectly classified as positive, resulting in an FP. Otherwise, it is considered a TN. The effectiveness of the attack is quantified using the FPR, allowing an assessment of the vulnerability of the classifier to adversarially generated inputs at inference time.
@@ 5.5 Flooding attack
A flooding attack is executed using a combination of FaN texts, paraphrased FaP, and rule-based generated FaP texts.
Paraphrasing TP texts. Multiple variations of a single real TP text are generated through paraphrasing. These texts serve as an effective means of flooding the model with slightly different yet semantically similar inputs. To achieve this, GPT-3.5-Turbo API was employed as a paraphrasing tool, applying it to the clean_tweet feature of the dataset. By generating ten paraphrases for each tweet from the positive class (1) 110740 FaP input texts were created.
@@ 5.6 Poisoning attack
To demonstrate the impact of a poisoning attack, FaN texts that were misclassified as positive (FP) by the model [14] and human analysts in the experimental evaluation of the evasion attack were gradually injected into the model training dataset over several retraining rounds. In each round, a random number of FaNs were injected to simulate realistic and unpredictable successive poisoning conditions. As a result, the classifier begins to gradually learn incorrectly, which impacts its classification decision boundaries.
@@ 5.7 Evaluation methodology
The attacks studied are characterized in Table II in terms of the adversarial text input type, purpose, and the corresponding classifier responses, explaining how the different attacks exploit the model’s vulnerabilities and the evaluation metrics used in the experiments. In evasion and flooding attacks, the adversary generates FaN texts to cause the model to misclassify them as FP samples, thus increasing the FPR. In flooding attacks, the adversary further intensifies the disruption by generating FaP texts, creating a flood of adversarial texts to overwhelm the system and deceptively increase the True Positive Rate (TPR). Moreover, the presence of FaN texts in flooding attacks amplifies evasion attacks. The poisoning attack can cause the model to misclassify input into FP or FN, which are better captured by the $\text{F}_{1}$ score.
@@
{tblr} cells = c, cell11 = r=2, cell12 = c=2, cell14 = c=3, cell34 = fg=red, cell41 = r=3, cell44 = fg=red, cell53 = r=2, cell54 = fg=red, cell56 = r=2, cell64 = fg=red, vlines, hline1,3-4,7,8 = -, hline2,5 = 2-6, hline6 = 2,4-5, Attacks & Adversary Model prediction
Generated fake text Intent Positive Negative Eval
Evasion FaN Misclassify as TP FP TN FPR
Flooding FaN Flooding dashboard, and
possibly amplify a FP evasion attack FP TN FPR
Paraphrasing, FaP Flooding the dashboard, amplify TP TP FN TPR
Rule-Based, FaP TP FN
Poisoning FaN Degrade decision boundaries FP FN $\text{F}_{1}$ score
@@ 6 Experimental results
where $A$ and $B$ represent the feature vectors of two tweets and $\mathbf{m}$ and $\mathbf{n}$ correspond to their respective embeddings. To visualize how the similarity scores are distributed within the dataset, Figure 5 presents a scatter diagram of all computed cosine similarity values between FaN texts and real positive tweets, sorted in ascending order. Each point reflects the semantic similarity between a generated FaN and a real positive tweet. The distribution shows that most scores fall between 0.5 and 0.8, indicating moderate semantic closeness. A smaller number of samples have scores below 0.5, while a few approach 1.0, suggesting high similarity in certain cases.
@@ 6.1 Evasion attack
False Positive Rate. The FPR measures how well the generated texts misled the classifier, representing the proportion of non-security generated texts incorrectly classified by the model as security-related. Table III compares the FPR achieved by ChatGPT-4o and the binary classifier proposed by [14], both evaluated on the original generated text and on the final text optimized by prompt refinement. ChatGPT-4o has a dual role in this experiment: as a generator of adversarial texts and a classifier responsible for detecting security-related content. This reveals a key weakness of the model, as it produced misleading samples but failed to classify them correctly. The results show a significant improvement in FPR due to prompt refinement. Regarding the models, the Dionisio et al. [14] model is significantly more vulnerable to adversarial attacks, with an FPR of 87% for the original texts and 97% for the optimized texts. Even with a very high FPR, ChatGPT-4o is significantly less vulnerable than the specialized model. These findings underscore how adversarial optimized prompt amplifies the FPR and reveal different levels of classifier robustness.
TABLE III: FPR (%) of adversarial text generated by ChatGPT-4o on the Twitter dataset [14], evaluated both in its original form and after optimization.
Models Original text Optimised text
ChatGPT-4o 0.51 0.75
[14] 0.87 0.97
[…]
@@
TABLE IV: Statistical comparison of gradients for real and FaN texts.
Metrics fake Real
Mean -1.10e-06 -1.28e-06
Variance 2.30e-05 2.67e-05
Wasserstein distance - 0.0012
KL divergence - 0.05
Cosine distance - 0.85
@@
Figure 6: Kernel Density Estimation Plot of real tweet and FaN text gradients.

### v2

@@ 5.2 Target models
Two models are targeted in the experiments. The first is a state-of-the-art binary classifier developed by Dionisio et al. [21] based on Convolutional Neural Networks (CNNs), employed within CTI extraction pipelines to recognize whether a tweet is related to cybersecurity or not. Binary labels and classifiers are fundamental requirements for the evasion attack considered in this study, as the generation of FaN texts explicitly relies on a binary decision boundary between security-related and non-security-related inputs. A successful evasion attack is defined as forcing a non-security text to cross this boundary and be misclassified as security-related, which cannot be unambiguously defined in multi-class or open-set settings.
We target the state-of-the-art binary classifier proposed by Dionisio et al. [21], which is based on Convolutional Neural Networks (CNNs) and used in CTI extraction pipelines to distinguish cybersecurity-related tweets from others. This choice is motivated by its recognised effectiveness [7] and training on real-world tweets, ensuring a realistic adversarial target. Moreover, the binary nature is essential for reliably defining evasion success in FaN text generation, because multi-class or open-set settings lack a clear decision boundary. The second target model is Chat( GPT-4o and GPT-5-mini), selected for exceptional performance in classifying security-related and non-security texts [59]. Unlike traditional classifiers, ChatGPT integrates contextual understanding with classification tasks to effectively detect subtle cyber threat patterns. Moreover, the model’s widespread adoption and versatility make it a practical benchmark for testing adversarial robustness in real-world scenarios. By targeting ChatGPT, we aim to evaluate the resilience of cutting-edge language models against evasion attacks. The evasion attack targets both models, while the flooding and poisoning attacks are specific to the specialized classifier [21].
@@ 5.3 Adversarial text generation
ChatGPT-4o was leveraged to generate texts that mimic the structure and terminology of cybersecurity texts while remaining unrelated to real cybersecurity. All adversarial texts were generated using ChatGPT-4o and ChatGPT-5-mini, each with a fixed prompt that preserves the lexical appearance and structural patterns of cybersecurity texts while explicitly removing security-related semantics. To ensure reproducibility, we used fixed generation parameters with a temperature of 0.7 and a maximum length limited to 100 tokens. While LLM-based text generation has limitations, such as potential biases, repetition, and inaccuracies [15], the method proposed in this study exceeded the expectations. To generate effective adversarial texts with ChatGPT-4o, prompts were designed using real tweets. Prompt optimization was performed iteratively. The initial prompt simply instructed the model to rewrite a cybersecurity-related text into a non-security text while preserving its surface structure and excluding security-related semantics. However, early outputs often leaked security semantics or appeared linguistically generic. Over five refinement iterations, we progressively introduced explicit constraints to: (i) forbid any security-related content; (ii) enforce transformation of key terms into visually similar but non-security tokens; (iii) preserve the original sentence structure; (iv) emphasize stylistic resemblance to CTI messages; and (v) encourage lexical diversity during generation. The final prompt, shown in Figure A1 in Appendix A, reflects these refinements and produces texts that are non-security in content while remaining deceptively similar to cybersecurity messages in their surface structure, thereby reducing semantic leakage and improving stylistic deception. An attention-based mechanism is proposed to extract the contextually key tokens from the real cybersecurity tweets. In the following, we first define the problem of generating adversarial text and then explain the generation process and its components as shown in Figure 3.
@@
Figure 3 illustrates the FaN text generation methodology [^2] Code will be available on https://github.com/samanehshf/Fake-CTIs. The attention mechanism identifies the top three tokens. In our experiments, the number of key tokens extracted using the attention mechanism was fixed to three. This choice strikes a balance between preserving sufficient cybersecurity-related lexical cues and minimizing excessive semantic leakage from the original content. Using fewer tokens (e.g., two) often resulted in adversarial texts that lacked sufficient stylistic resemblance to cybersecurity texts, whereas using more tokens (e.g., five) increased the risk of retaining security-specific semantics. To validate this design choice, we conducted a series of smaller-scale experiments in which we varied the number of key tokens and evaluated their impact on prompt effectiveness. Across all settings, using three tokens led to a better deceptive effect and achieved the highest evasion performance. Then ChatGPT-4o modifies the surrounding top three tokens as suggested in the prompt to create adversarial texts that still appear cybersecurity-related. The prompt follows an iterative refinement, where each cycle evaluates the generated FaNs based on their ability to deceive the pre-trained text classifier used in our pipeline. The prompt is considered adequate if the classifier produces a number of FPs (#FP\#FP) exceeding a predefined threshold $\vartheta$. Otherwise, another refinement iteration is executed. The finalized prompt was effective enough to surpass the $\vartheta$ setting of 80% of the input texts. It was then used to generate the adversarial FaN texts, forming the basis for subsequent evaluation.
Human evaluation. A human evaluation step was included after the final prompt optimization iteration to further validate the effectiveness of the generated FaN texts. To filter out low-quality fakes, the data were distributed to three analysts who were informed that the texts were fake and tasked with identifying those that resembled cybersecurity texts. The analysts had prior experience in cybersecurity-related research and threat intelligence analysis, and were asked to base their judgments solely on linguistic plausibility, domain-specific terminology, and overall resemblance to real-world CTI texts. These three dimensions operationalize the notion of “resemblance” and provide a consistent basis for human judgment. Analysts were clearly informed that the texts were fake to prevent external fact-checking or online verification, thereby enforcing a conservative evaluation setting in which texts had to appear realistic even under careful review. This result was crucial to achieving a final prompt that surpassed the $\vartheta$ threshold. The same analysts were later involved in downstream validation stages of the CTI pipeline, reflecting a realistic analyst-in-the-loop workflow and avoiding redundant human evaluations. The importance of the iterative prompt refinement with human validation is demonstrated by the difference between the initial and final FPR presented in Section 6.
@@ 5.4 Evasion attack
The evasion attack was conducted by passing the adversarially generated FaN texts through the target classifiers. A successful attack occurs when a FaN is incorrectly classified as positive, resulting in an FP. Otherwise, it is considered a TN. The effectiveness of the attack is quantified using the FPR, allowing an assessment of the vulnerability of the classifier to adversarially generated inputs at inference time. In contrast to classical adversarial NLP attacks that rely on constrained token-level perturbations or explicit optimization of a loss function, the evasion attack considered in this work is generation-based. It models a realistic CTI threat scenario in which adversaries inject fully fake cybersecurity-like texts into OSINT sources, rather than minimally modifying existing samples. This design choice reflects real CTI pipelines, where attackers freely generate deceptive content at scale using public LLMs without edit budgets or semantic preservation constraints.
@@ 5.5 Flooding attack
The flooding attack is executed using a combination of FaN texts, paraphrased FaP, and rule-based generated FaP texts, to study the efficacy of different generation methods and text types.
Paraphrasing TP texts. Multiple variations of a single real TP text are generated through paraphrasing. These texts serve as an effective means of flooding the model with slightly different yet semantically similar inputs. To achieve this, GPT-3.5-Turbo API was employed as a paraphrasing tool, applying it to the clean_tweet feature of the dataset. By generating ten paraphrases for each tweet from the positive class (1), 11074 FaP input texts were created.
@@ 5.6 Poisoning attack
To demonstrate the impact of the poisoning attack, 9402 FaN texts that were misclassified as positive (FP) by the model [21] and human analysts in the experimental evaluation of the evasion attack were gradually injected into the model training dataset over several retraining rounds. In each round, a random number of FaNs were uniformly sampled and removed from the 9402 set, and injected into the training data to simulate realistic and unpredictable successive poisoning conditions. As a result, the classifier gradually begins to learn incorrectly as the cumulative number of FaN entries grows, affecting its classification decision boundaries and performance.
@@ 5.7 Evaluation methodology
The attacks studied are characterized in Table II in terms of the adversarial text input type, purpose, and the corresponding classifier responses, explaining how the different attacks exploit the model’s vulnerabilities and the evaluation metrics used in the experiments. In evasion and flooding attacks, the adversary generates FaN texts to cause the model to misclassify them as FP samples, thus increasing the FPR. In this context, the FPR directly corresponds to the attack success rate (ASR), as a successful attack is defined by the misclassification of adversarial inputs as security-related. In flooding attacks, the adversary further intensifies the disruption by generating FaP texts, creating a flood of adversarial texts to overwhelm the system and deceptively increase the True Positive Rate (TPR). Moreover, the presence of FaN texts in flooding attacks amplifies evasion attacks. The poisoning attack can cause the model to misclassify input into FP or FN, which are better captured by the $\text{F}_{1}$ score.
@@
Attacks Adversary Model prediction Generated fake text Intent Positive Negative Eval Evasion FaN Misclassify as TP FP TN FPR Flooding FaN Flooding dashboard, and possibly amplify a FP evasion attack FP TN FPR Paraphrasing, FaP Flooding the dashboard, amplify TP TP FN TPR Rule-Based, FaP TP FN Poisoning FaN Degrade decision boundaries FP FN $\text{F}_{1}$ score
5.8 Attack cost
The attack budgets are characterized to assess their real-world feasibility and reproducibility. All attacks rely on standard usage of publicly accessible LLM APIs. For the evasion attack, the cost scales linearly with the number of adversarial texts generated. Each FaN sample is generated with a single LLM query, using a fixed prompt without fine-tuning. This keeps the evasion attack lightweight and feasible under a modest query budget. Flooding attacks require minimal additional cost beyond text generation because adversarial samples can be generated offline and injected into the CTI pipeline. The main constraint is the CTI pipeline’s ingestion rate, rather than a moderate computational complexity, as large volumes of adversarial text can be generated using the same prompt template and injected into the system. Poisoning attacks are the most resource-intensive, as they require continuous attacker effort over a prolonged period to generate a diverse set of fake texts that can evade the classifier, reach the model’s training data, and change its decision boundary. The results shown in Table VI confirm the higher cost, showing that the need to inject a few thousand misclassified samples over successive retraining rounds is required to substantially degrade model performance.
@@ 6 Experimental results
where $A$ and $B$ represent the feature vectors of two tweets and $\mathbf{m}$ and $\mathbf{n}$ correspond to their respective embeddings. To visualize how the similarity scores are distributed within the dataset, Figure 5 presents a scatter diagram of all computed cosine similarity values between FaN texts and real positive tweets, sorted in ascending order. Each point reflects the semantic similarity between a generated FaN and a real positive tweet. The distribution shows that most scores fall between 0.5 and 0.8, indicating moderate semantic closeness. The dashed horizontal lines indicate heuristic similarity thresholds (50% and 80%) used to categorize tweets into low, moderate, and high semantic similarity groups. A Shapiro–Wilk [60] normality test indicates that the similarity distribution does not follow a normal distribution ($p=3.2\times 10^{-6}$). This result, along with a mean of 0.62, indicates a positive skew in the distribution, consistent with the idea that most texts exhibit considerable similarity.
@@ 6.1 Evasion attack
False Positive Rate. The FPR measures how well the generated texts misled the classifier, representing the proportion of non-security generated texts incorrectly classified by the model as security-related. Table III compares the FPR achieved by ChatGPT-4o and the binary classifier proposed by [21], both evaluated on the original generated text and on the final text optimized by prompt refinement. ChatGPT-4o has a dual role in this experiment: as a generator of adversarial texts and a classifier responsible for detecting security-related content. This reveals a key weakness of the model, as it produced misleading samples but failed to classify them correctly. The results show a significant improvement in FPR due to prompt refinement.
To assess the stability of the results, the optimized evasion attack was repeated three times with fixed prompts and LLM parameters. Table III reports the mean FPR and standard deviation across three runs for both Twitter and AnnoCTR datasets. We additionally report the $t$-interval 95% confidence interval (CI) for each optimized experiment, to quantify the uncertainty induced by the stochastic nature of LLM-based text generation.
Regarding the Twitter data set, the Dionisio et al. [21] model is significantly more vulnerable to adversarial attacks, with an FPR of 87% for the original texts and 97% for the optimized texts. Even with a very high FPR, ChatGPT models are significantly less vulnerable than the specialized model.
The same evaluation protocol was applied to the AnnoCTR corpus. Table III reports the results on the full set of sampled AnnoCTR text segments and on the smaller subset of longer texts. These ranged from 256 to 1000 characters and reflect the behavior of the models under more realistic CTI inputs. Comparing the results, it may be observed that, in general, the evasion attack is more effective against ChatGPT models on the full AnnoCTR text set than against the Twitter dataset. The effectiveness is equal when considering the specialized model. Interestingly, the results on the subset of longer AnnoCTR texts are similar across all models to those on the Twitter dataset, but less effective for the ChatGPT models compared to the full AnnoCTR texts. Globally, within the range of text size considered in both datasets (from less than 256 to 1000), the results are not significantly different. The most significant difference is made apparent between the specialized model and the ChatGPT models, with an advantage to ChatGPT-4o.
Considering the success of attacks, the human evaluation, the cosine similarity results, and a visual inspection on a sample of adversarial texts, from a linguistic perspective, the misclassified adversarial samples closely mimic the surface structure of genuine CTI messages. They preserve domain-like terminology, identifier-shaped tokens resembling CVE or software component references, and CTI-style sentence structures, despite containing no actual security semantics. Such results indicate that the target classifiers predominantly rely on shallow lexical and structural cues, rather than assessing the underlying semantic validity of the content.
TABLE III: FPR (%) of adversarial text generated on the Twitter [21] and AnnoCTR datasets. Results for optimized text are reported as mean $\pm$ standard deviation over three runs. For each optimised column, we additionally report the 95% confidence interval (CI) computed using a $t$-interval.
Twitter AnnoCTR AnnoCTR (longer texts) Models Original text Optimised text (mean $\pm$ std, 95% CI) Optimised text (mean $\pm$ std, 95% CI) Optimised text (mean $\pm$ std, 95% CI) ChatGPT-5-mini $51.0$ $68.0\pm 1.2$ $[65.0,71.0]$ $74.0\pm 2.5$ $[67.8,80.2]$ $69.3\pm 1.5$ $[65.6,73.0]$ ChatGPT-4o $51.0$ $75.0\pm 1.7$ $[70.8,79.2]$ $81.0\pm 1.7$ $[76.8,85.2]$ $76.3\pm 1.5$ $[72.6,80.0]$ Dionisio et al.[21] $87.0$ $97.0$ $97.0$ $96.0$
Moreover, to complement the fixed-threshold FPR analysis, in Figure 6 we present receiver operating characteristic (ROC) and precision–recall curves for the Dionisio et al. [21] model under evasion attack. These threshold-independent evaluations provide additional insights into the classifier’s robustness to adversarial FaN samples. The evaluation was conducted on a mixed test set consisting of 11074 real cybersecurity-related texts, 5000 non-security texts, and 5000 adversarially generated FaN texts designed to resemble cybersecurity content. The area under the curve (AUC) is 0.523, which is close to random guessing, indicating that the classifier fails to distinguish between real security-related texts and adversarially generated non-security texts across various decision thresholds. The near-diagonal ROC curve confirms that evasion attacks significantly degrade the model’s discriminative capability at inference time. Figure 6b illustrates the Precision–Recall (PR) curve, with an average precision (AP) of 0.533. Despite maintaining moderate recall, precision remains low across most thresholds, reflecting a high FPR caused by FaN texts. This indicates a reliance on surface-level lexical and structural patterns instead of robust semantic verification, leaving the classifier open to exploitation by generative adversarial inputs.
(a) ROC curve
(b) Precision–Recall curve
Figure 6: Performance of the [21] model under evasion attack across decision thresholds. (A) Receiver operating characteristic (ROC) curve showing near-random discrimination. (B) Precision–recall curve indicating a high false alarm rate.
Gradient distribution. Gradients can be used to quantify the model’s sensitivity to minor input variations, being essential for assessing robustness against adversarial attacks [24]. The target binary classifier model [21] is assessed by evaluating FaN adversarial text gradient behavior against the baseline gradient of real texts. Kernel density estimation [32] is employed, which is a non-parametric technique for estimating the Probability Density Functions (PDFs). Figure 7 (a) and (b) present PDF plots of gradient distributions for real (blue) and adversarial (yellow) samples across datasets. Dashed vertical lines mark the mean of each distribution. Subfigures (c) and (d) present relevant details for comparing the dataset distributions of real (c) and FaN texts (d).
For both datasets, the real and FaN text distributions exhibit similar Gaussian-like shapes with near-zero means. This demonstrates three key points: (1) adversarial inputs closely mimic the overall gradient structure of real texts, (2) subtle differences (narrower variance in adversarial gradients) offer potential for attack detection, (3) some high-magnitude gradients in real texts are replaced by near-zero values, indicating feature cancellation.
Subfigures (a) and (b) induce the same conclusions for the Twitter and AnnoCTR datasets: real CTI reports and generated FaN texts have highly overlapping distributions with only minor differences in peak height and concentration. Subfigures  (c) and  (d) compare the datasets regarding real and FaN text. Real samples (c) show nearly identical gradient distributions between Twitter and AnnoCTR, indicating the stability of the data sources. However, fake samples (d) exhibited somewhat greater variations in peak height and concentration, reflecting the data source’s sensitivity to adversarial generation. Supporting metrics, including gradient means, variances, and distributional distances, are reported in Table IV for both datasets. The adversarial gradient mean and variance closely match real text, suggesting effective imitation. However, the slightly narrower variance of adversarial gradients indicates reduced content diversity, likely due to constraints in the adversarial generation process. The KL divergence ($\text{fake}\rightarrow\text{real}$) confirms that both PDFs are very similar, reinforcing the effectiveness of the attack. The cosine distance confirms the directional similarity between gradients, indicating that adversarial inputs affect the model while preserving detectable features. Finally, the Wasserstein distance [52] suggests a strong geometric alignment between adversarial and real gradients [43].
@@
TABLE IV: Statistical comparison of gradient distributions between real and FaN texts on the Twitter and AnnoCTR datasets.
Twitter AnnoCTR Metrics fake Real fake Real Gradient mean -1.10e-06 -1.28e-06 -0.90e-06 -1.4e-06 Gradient variance 2.30e-05 2.67e-05 2.0e-05 2.5e-05 Wasserstein distance 0.0012 0.0007 KL divergence 0.05 0.06 Cosine distance 0.85 0.88
@@
(a) Twitter: Real vs FaN
(b) AnnoCTR: Real vs FaN
(c) Real: Twitter vs AnnoCTR
(d) Fake: Twitter vs AnnoCTR
Figure 7: KDE of gradient distributions across datasets.